OpenStudy (anonymous):

Solve using the quadratic formula: x = -(-14)+- sqrt(-(-14)^2=4(-2)(12)/2(-2)

OpenStudy (anonymous):

so far I have 14 +- sqrt(196+96)/-4 = 14+- sqrt(292)/-4

OpenStudy (anonymous):

-2x^2-14x+12? is that the original problem?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{ -14\left(\begin{matrix}+ \\ -\end{matrix}\right)\sqrt{(-14)^2-4(-2)(12)} }{ 2(-2) }\]

OpenStudy (anonymous):

my bad -14 in the beginning should be +14

OpenStudy (anonymous):

I got that far but going from there is my problem

OpenStudy (anonymous):

don't think 292 is right

OpenStudy (anonymous):

it is

OpenStudy (anonymous):

so then you have 14 +- sqrt(292) / -4 and how do you simplify that? to a decimal?

OpenStudy (anonymous):

check if the original problem is correct, you might have written it down wrong

OpenStudy (anonymous):

the original problem is -2x^2 - 14x + 12

OpenStudy (anonymous):

@eHoaX

OpenStudy (anonymous):

ok so everything is correct

OpenStudy (anonymous):

I am finding the x-intercepts so I can graph the parabola

OpenStudy (anonymous):

but when I put it into wolfram alpha I get a decimal of 9.7...

OpenStudy (anonymous):

it should be .7 and -7.7

OpenStudy (anonymous):

not sure what I am doing wrong

OpenStudy (anonymous):

i am also getting -7.77 and .77

OpenStudy (anonymous):